# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v3

@@ 6.2 Mapping Pillars of TRiSM with Agentic AI
Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art. 13–14) [106] and the NIST AI RMF “Explain” function [107]. This pillar promotes transparent, rationale, human-interpretable logging, and role-based interpretability [108]. ModelOps : Anchored in the NIST RMF “Govern” function [107] and ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]. It emphasizes lifecycle controls, model versioning, rollback policies, and CI/CD pipelines with safety gates. Application Security: Guided by OWASP LLM Top-10 [102] and OECD AI Principles on Robustness [109], it covers access controls, prompt sanitization, sandboxing, and audit trail mechanisms. Model Privacy: Aligned with GDPR (Art. 25) [103], HIPAA §164 [104], and OECD AI Principles [109], this pillar ensures data minimization, encryption, differential privacy budgeting, and access-audit pass thresholds. Governance: Grounded in overarching regulatory frameworks such as GDPR, CCPA [110], and the OECD AI Principles [109], this pillar supports institutional oversight, policy compliance, and cross-system accountability. It enforces auditable workflows, verifiable role-based access, regulatory traceability, and human oversight, particularly in multi-agent and high-risk AI deployments.
Figure 6: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST RMF, ISO/IEC 42001, 42005, and OECD Principles. Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
Explainability in Multi-Agent Decision Making
Explainability refers to making the inner workings and decisions of AI agents interpretable to humans. In the context of LLM-based multi-agent systems, it is paramount for building user trust because outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction [111]. However, achieving explainability in such systems is particularly challenging , one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to the final outcomes [112]. In other words, explaining why a multi-agent system behaved a certain way requires tracing which agent did what and how their collaboration produced an emergent result, a task that traditional single-model explainability techniques do not directly address. Several approaches are being explored to enhance explainability in LLM-driven multi-agent systems:
  • Feature Attribution (LIME/ SHAP): Established explainable AI techniques provide a starting point. For instance, LIME) [113] and Shapley Additive Explanations (SHAP) [114] can be adapted to analyze LLM decisions. These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken. In an Agentic LLM setting, a “feature” might be a component of the agent’s input context or a signal from another agent. By highlighting the most influential parts of each agent’s input, methods like LIME/SHAP help clarify individual agent decisions in the multi-agent mix [115].
@@
  • Counterfactual Analysis Beyond local feature attributions, “what-if” analyses are increasingly important for multi-agent explainability. Counterfactual techniques examine how the system’s behavior would change if certain conditions or agent behaviors were altered [116]. For example, one can remove or modify a particular agent’s contribution and then observe how the collective outcome changes. This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making.
@@
  • Reasoning Trace Logging: Instead of treating the agents as black boxes, an explicit trace of their intermediate reasoning can greatly enhance transparency. Logging the step-by-step reasoning process – for example, an agent’s chain-of-thought (CoT) prompts or the dialogue exchanged between agents – creates an audit trail of how the decision emerged [69]. By examining this trace, one can see how agents influence each other, which agent contributed what, and how partial conclusions build up to the final decision. This visibility into the multi-agent process addresses a key blind spot of standard XAI, effectively “opening up” the black box of inter-agent interactions.
  • Natural Language Explanations: Finally, the generative nature of LLMs can be leveraged to turn the above insights and logs into human-friendly explanations [117]. Instead of expecting end-users to read raw CoT logs or analyze feature weights, an “explainer” agent (which could be another LLM) or a human analyst can compile the reasoning trace into a narrative. In fact, an important aspect of explainability is communicating the reasons behind decisions in an intelligible way. By using natural language, the system can answer user questions like “Why did the agents decide to do X?” in a conversational manner.
ModelOps : Lifecycle Management for LLM Agents
ModelOps (Machine Learning Operations) is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement [118]. Within AMAS, ModelOps covers not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely [59]. Effective ModelOps helps maintain consistency, performance, and regulatory compliance as LLM agents evolve or are added to a system. AMAS require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental variations. A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations. Additionally, robust CI/CD pipelines (Continuous Integration/Continuous Deployment) are employed to automatically test agents’ performance and safety whenever a model is fine-tuned or an agent’s logic is modified. Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions.
Application Security for LLM Agents
To mitigate threats such as prompt injection, spoofing, and impersonation, the TRiSM security pillar recommends a defense-in-depth approach tailored to LLM agents. This includes prompt hygiene [19] , where agents treat inputs defensively by sanitizing and filtering prompts, and applying guardrails or content policies to detect and refuse suspicious instructions. Techniques like prompt hardening, such as secure prefixes or validation steps, can help reduce susceptibility to injection attacks [119]. It also involves strong authentication and access controls, ensuring that each agent and human user is securely authenticated, and that least-privilege principles are enforced to limit autonomous agent actions [120]. Finally, continuous monitoring plays a critical role: if an agent begins issuing anomalous requests or deviates from typical behavior patterns, automated systems can flag the incident for review or trigger a shutdown to prevent further risk.
@@
Recent frameworks, such as LangChain [121], AutoGen [122], and CrewAI [81] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise. Moreover, training LLM agents with adversarial robustness in mind (e.g. fine-tuning on adversarial examples, employing adversarial training regimes) [19] can improve their resistance to malicious inputs.
Model Privacy and Data Protection
The model privacy addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms [21]. Since multiple agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate. Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs. To preserve privacy in AMAS, several technical measures can be employed. For example, Differential Privacy (DP) adds calibrated noise during model training to ensure that no individual data point can significantly influence the output, thus preventing memorization of personal information [123]. Data anonymization and minimization further enhance privacy by limiting shared information to only what is essential, often in aggregated or pseudonymized formats, such as using “age bracket 30–40” instead of precise birthdates [124]. Secure Multi-Party Computation (SMPC) allows agents to jointly compute functions without revealing their private inputs, making it particularly useful in cross-organizational scenarios like collaborative fraud detection [125]. Similarly, Homomorphic Encryption (HE) permits computation directly on encrypted data; with Fully Homomorphic Encryption (FHE), even plaintext queries and responses remain inaccessible to the agent itself [126].Trusted Execution Environments (TEEs), or secure enclaves, offer hardware-based isolation, ensuring that even privileged users cannot access data during processing, which is an essential safeguard for secure memory sharing and execution [127].
Governance
Governance of AI systems necessitates compliance with evolving legal and policy frameworks, including the General Data Protection Regulation (GDPR) [103] and the California Consumer Privacy Act (CCPA) [110]. Effective governance extends beyond regulatory adherence to cover proactive management of legal, ethical, and societal risks. This includes establishing transparent accountability structures, maintaining auditable documentation, and ensuring the protection of data subject rights throughout the AI lifecycle.
Several global frameworks provide guiding principles that align with the TRiSM pillars. The EU AI Act [106], particularly Article 14, mandates transparency, traceability, and human oversight in high-risk systems; principles embedded within TRiSM’s Explainability and ModelOps dimensions. Similarly, the NIST AI Risk Management Framework (AI RMF) [107] identifies: Explainability, Governability, and Validity as key components of AI risk mitigation, which directly correspond to TRiSM’s lifecycle control mechanisms. The OECD AI Principles [109] reinforce the importance of robustness, transparency, and accountability, echoing TRiSM’s attention to secure tool invocation, memory traceability, and multi-agent auditability.
Together, these frameworks underscore the need for governance architectures in Agentic AI systems that support real-time monitoring, role-based access, verifiable logs, and intelligible outputs. As illustrated in Figure 6, TRiSM offers a practical scaffolding for operationalizing these requirements in LLM-based multi-agent environments.
7 Trust and Explainability in Agentic AI Systems
Table 5: Trust-Enabling Mechanisms in Agentic AI Systems
Mechanism Purpose Example System(s) Trust Contribution
Transparency via Reasoning Traceability Make agent decisions observable through intermediate steps or explanations SciAgent [128], MetaGPT [129] Enhances interpretability and debuggability
Status Reporting and Intent Disclosure Provide real-time updates or clarification of progress and errors ChatDev [13], AutoGen [122] Mimics human team transparency; reduces uncertainty
Human-in-the-Loop (HITL) Enable user oversight over high-stakes or irreversible actions ChemCrow [130], enterprise dashboards Prevents unsafe execution; reinforces control
Behavioral Consistency & Bounded Autonomy Ensure agents operate within predefined roles and limits Adaptive tutors [131, 132] Improves predictability; prevents overreach
Social Trust Cues Use polite language, uncertainty and expressions GPT-4 safety-tuned profiles [133, 134] Enhances user comfort and trust in uncertain scenarios
Two key factors that influence trust in Agentic AI are the transparency of the agent’s decision-making processes and the ability to explain or justify its actions in human-understandable terms. Figure 7 illustrates the layered relationship between trust mechanisms, explainability techniques, and evaluation facets. Trust-enabling mechanisms, such as transparency and human-in-the-loop (HIIL) oversight, provide the foundation for user confidence. These feed into explainability techniques like decision provenance graphs, which make agent behaviors interpretable. Finally, evaluation (Section 8) facets ensure that trust and explainability are quantifiable, enabling iterative improvements in system design.
7.1 Building Trust in Autonomous Agents
In Agentic AI systems, building trust is foundational to user acceptance, system reliability, and responsible deployment, especially as these systems begin to make autonomous decisions in critical domains such as healthcare, finance, and scientific research [135]. Unlike traditional software agents, autonomous LLM-based agents are characterized by self-directed reasoning, adaptive memory, and dynamic collaboration, which make their operations opaque and often unpredictable [136]. Establishing trust in such systems, therefore, requires a combination of technical transparency, user feedback integration, and robust oversight mechanisms.
Trust-Enabling Mechanisms Transparency • Status Reporting • HITL • Bounded Autonomy • Social Cues Explainability Techniques LIME • SHAP • Decision Provenance • Attention Maps • Prompt Audits Evaluation Facets Trustworthiness • Explainability • User-Centric • Coordination • Composite
Figure 7: Relationship between trust mechanisms, explainability techniques, and evaluation facets in Agentic AI.
Transparency and Decision Traceability. Transparency is one of the core enablers of trust. For users to understand and evaluate AMAS decisions; the reasoning chains, decision states, and action triggers of agents must be made observable. Several Agentic AI systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules. For example, SciAgent [128] generates scientific summaries and provides justifications by linking outputs to source documents retrieved via retrieval-augmented generation (RAG). Similarly, MetaGPT[129] structures its reasoning using role-based outputs, where each agent (e.g., planner, coder) explicitly states the logic behind its task execution, creating modular interpretability.
Status Reporting and Progress Visibility. Clear reporting of intent and intermediate status is also essential. Human collaborators often require updates about what an agent is doing, why a task is taking longer, or how an agent interprets ambiguous instructions. Tools like AutoGen[122] and ChatDev[13] have incorporated structured chat interfaces where agents summarize their intermediate progress, decisions, and encountered errors.
Human-in-the-Loop Oversight. Human oversight and intervention mechanisms further reinforce trust. Allowing human users to review, edit, or approve agent-generated outputs not only prevents missteps but also signals that the system respects user authority [137, 138, 139]. Many systems adopt a human-in-the-loop (HITL) paradigm [140], where agents request confirmation before executing high-risk or irreversible actions.
Behavioral Consistency and Bounded Autonomy. Trust requires predictability in models’ behaviour. Agents should follow defined roles, output in expected formats, and remain within delegated authority. For example, in enterprise AI platforms used for automated data analytics, agents may generate insights or dashboards but defer publishing until a domain expert reviews the material. Similarly, in autonomous research assistants like ChemCrow [130], agents pause to allow chemists to validate proposed reactions or data pipelines before proceeding, reinforcing safe deployment.
Social Trust Cues and Language Behaviors. Beyond system-level mechanisms, behavioral consistency and guided autonomy are crucial. If an agent behaves unpredictably or inconsistently, even if technically correct, users are less likely to trust it. Behavioral alignment mechanisms such as predefined role protocols can help output style consistency [141, 12, 142]. For example, in adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries . Also, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings [143]. Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence [134]. This has been implemented in models like GPT-4 when configured with safety-tuned instruction sets, improving reliability perception without undermining capability. These mechanisms form a layered trust strategy for Agentic AI, as summarized in Table 5.
7.2 Explainability Techniques in AMAS
Emergent behavior of AMAS poses a challenge: the interpretability of an individual agent does not necessarily imply the interpretability of the overall system. In platforms such as ChatDev [13] or AutoGen [53], agents emulate specialized roles (e.g., engineer, reviewer), and tracing a final action back to its source agent is often non-trivial. While local explainability techniques such as LIME [113] and SHAP [114] remain foundational, their application in Agentic AI is limited due to decentralized agent contexts and shared memory. Researchers have thus proposed composite frameworks that combine local attribution with global decision traceability, such as decision provenance graphs [144], causal influence chains [145], and multi-agent SHAP adaptations [146]. These techniques aim to surface the system-level dynamics that underlie emergent agent behaviors.
Symbolic and Hybrid Architectures. Another promising direction is the use of inherently interpretable modules such as rule-based planners and decision trees within hybrid architectures [5]. These agents offer built-in explainability but retain the generative capabilities of LLMs for broader context understanding. Such designs are increasingly used in domains where structure and interpretability are prioritized, such as educational AI or mission planning.
Lightweight Interpretability: Attention and Prompts. Attention map visualizations have been used to highlight focus areas in multi-modal language agents [147, 148, 149], offering lightweight but informative views into model behavior. Prompt audit trails logging prompt history, agent actions, and response metadata have also gained traction [150, 151]. These mechanisms support system debugging, safety evaluations, and human-in-the-loop fine-tuning in multi-agent environments.
Table 6: Explainability Techniques in Agentic AI Systems
Technique Purpose Challenges in Agentic AI Example Use Cases
LIME [113] Local post-hoc explanation using surrogate models Fails to capture cross-agent dependencies in multi-agent setups Feature-level auditing in fraud detection
SHAP [114] Attribution of prediction to input features via Shapley Difficult to aggregate across agents with divergent contexts Financial risk analysis, regulatory explanations
Decision Provenance Graphs [144, 152] Trace inter-agent decision paths and data flows Can be complex to interpret at scale Collaborative document generation, planning chains
Multi-agent SHAP [146] Attribution extended to agents and shared memory High complexity, lack of standard frameworks Strategic gameplay, collaborative writing
Symbolic/Rule-Based Agents [5] Built-in interpretability via logic or rule-based models Limited flexibility compared to LLMs Hybrid planning systems, educational tutors
Attention Maps [147, 149] Visualize focus of language/image models Only reflects internal focus, not logic or intent Multimodal VQA, image captioning
Prompt Audit Trails [150, 151] Log input prompts, outputs, and agent actions Requires infrastructure, may miss implicit state changes Debugging multi-agent sequences, fine-tuning
The integration of trust-enabling mechanisms and explainability techniques is critical for the adoption of Agentic AI systems in high-stakes domains. While transparency, human oversight, and interpretable architectures improve user confidence, challenges in capturing system-level dynamics and scaling explainability remain. Future research should focus on developing standardized frameworks for multi-agent explainability, real-time interactive querying, and longitudinal trust evaluation to ensure robust and responsible deployment. A comparative summary of current techniques is shown in Table 6.

### v4

@@ 5.2 Mapping Pillars of TRiSM with Agentic AI
In this section, we examine how the pillars of TRiSM align with Agentic AI (or AMAS) and summarize our syntheis in Table LABEL:tab:trism-pillars. For each pillar, we outline its relevance and implications within the context of Agentic AI.
5.2.1 Explainability in Multi-Agent Decision Making
Explainability refers to making the inner workings and decisions of AI systems interpretable to humans [145]. In AMAS settings, outcomes emerge from the interactions among specialized agents rather than a single model prediction [146]. Therefore, explainability must cover both individual agent decisions and the inter-agent dynamics that lead to emergent results. Recent research emphasises that user trust hinges on understanding how each agent contributes and how their collaboration is orchestrated [111]. Traditional single-model XAI techniques often ignore this multi-agent dimension, so new methods are required.
Challenges Explaining multi-agent decisions requires tracing which agent performed which sub-task, what information they used, and how their interactions influenced the final outcome [56]. Additional challenges include: (1) Emergent behaviour: interactions among agents can lead to behaviour that was not explicitly programmed; explaining these emergent patterns requires capturing interaction chains.; (2) Verification of intermediate steps: vanilla CoT prompting in LLMs often produces plausible-sounding but unchecked reasoning. Without verification, small errors can propagate
Recent directions Recent works introduce new frameworks and techniques for multi-agentic explainability:
  • Layered Chain-of-Thought prompting: Instead of producing a single monolithic reasoning trace, Layered-CoT divides reasoning into discrete layers [111]. Each layer answers a sub-question, and a specialized Reasoning Agent generates the partial chain-of-thought while Verification Agents and User-Interaction Agents cross-check the intermediate steps against knowledge graphs or human feedback. This agent-based decomposition localizes errors, prevents contradictions, and ensures that each partial solution is validated before the model proceeds.
  • Integrated multi-agent workflows: Multi-agent workflows embed explainability into the process itself rather than as a post-hoc add-on. For example, in a DSM-5 mental-health diagnosis workflow, one agent collects patient information while a diagnostician agent retrieves relevant DSM-5 criteria and explicitly links conversational utterances to diagnostic rules. This integration produces transparent, step-by-step diagnostic predictions and shows why specific recommendations are made [147]. The study notes that step-by-step reasoning improves perceptions of fairness and reliability and that embedding a diagnostician agent moves beyond general synthetic dialogues by providing a self-documenting rationale.
  • Modular resume-screening systems: In recruitment, a multi-agent architecture divides the process into extraction, evaluation, summarization and formatting. The evaluation agent dynamically retrieves company-specific hiring criteria via Retrieval-Augmented Generation (RAG), and each stage remains clearly defined. This modularity allows recruiters to trace how a candidate was assessed and why a particular score was assigned [148].
  • Explainable credit assignment in multi-agent reinforcement learning: A centralized LLM critic can provide individualized credit assignments to agents and explain its feedback strategy. Unlike earlier deep RL methods, this approach uses the language-based critic to justify why each agent receives a particular reward and to show how it decomposes the global reward [149].
Approaches for explainability in multi-agent LLM systems The following techniques adapt traditional XAI methods to the multi-agent setting:
  • Feature Attribution (LIME/ SHAP) for agents: Techniques like LIME [112] or SHAP [113] can be extended to attribute an agent’s output to specific input components (e.g., particular context tokens or inter-agent messages). Highlighting which parts of the input or which inter-agent signals most influenced an agent’s action helps illuminate why a specific agent behaved as it did [150].
@@ 5.2.1 Explainability in Multi-Agent Decision Making
  • Counterfactual and Causal Analysis Beyond local feature attributions, “what-if” analyses are increasingly important for multi-agent explainability. Counterfactual techniques examine how the system’s behavior would change if certain conditions or agent behaviors were altered [151]. For example, one can remove or modify a particular agent’s contribution and then observe how the collective outcome changes. This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making.
@@
  • Reasoning Trace Logging: Instead of treating the agents as black boxes, an explicit trace of their intermediate reasoning can greatly enhance transparency. Logging the step-by-step reasoning process – for example, an agent CoT prompts or the dialogue exchanged between agents – creates an audit trail of how the decision emerged [75]. By examining this trace, one can see how agents influence each other, which agent contributed what, and how partial conclusions build up to the final decision. This visibility into the multi-agent process addresses a key blind spot of standard XAI, effectively “opening up” the black box of inter-agent interactions.
  • Natural Language Explanations: Multi-agent systems can include a separate explainer agent that synthesizes logs and feature attributions into human-readable narratives. This agent (often another LLM) translates the technical reasoning traces into conversational explanations that answer questions like “Why did the agents choose action X?” Integrated workflows, such as DSM5AgentFlow [147], demonstrate that embedding such explanation directly in the pipeline improves transparency and trust.
  • Role-oriented modular architectures: Assigning clear roles to agents (e.g., reasoning, verification, summarization, credit assignment) enhances explainability because each stage of decision-making is distinct and can be inspected independently [148]. Modular designs also allow domain experts to adjust or replace specific agents without affecting the entire system.
  • Explainable reward critics: In multi-agent reinforcement learning, LLM-based critics can decompose global rewards and provide natural-language justifications, making credit assignment transparent [149] . This approach helps developers understand how credit is distributed across agents and why certain behaviors are encouraged or discouraged.
5.2.2 ModelOps : Lifecycle Management for LLM Agents
ModelOps (Machine Learning Operations) extends the MLOps paradigm beyond model training and deployment to include governance, risk management and continuous oversight for all types of AI models [131]. Unlike MLOps, ModelOps manages not only individual models but also composite systems of agents and heterogeneous models across different business units.
ModelOps provides end-to-end oversight for AI in production by enforcing model governance and compliance: tracking lineage, ownership, and metadata to meet regulations such as GDPR and HIPAA , and by using policy-driven deployment gates so only approved models go live, with controlled roll-outs and roll-backs [132]. It continuously monitors risk and bias, detecting drift, fairness issues, hallucinations, and robustness regressions. ModelOps also manages heterogeneous assets: rule-based systems, LLMs, and composite multi-agent architectures; ensuring consistent visibility and lifecycle control across the portfolio. Finally, it integrates tightly with enterprise IT, wiring AI workflows into CI/CD, IT service management (ITSM), identity and access management (IAM), and cloud-security tooling [132]. For generative AI and AMAS, ModelOps must handle prompt versioning, guardrails and content filtering enforcement, fine-tuning lineage tracking, usage and cost observability, and ethical compliance. Effective ModelOps thus requires:
  • Version control and CI/CD pipelines: Track versions of each agent’s model and prompt configuration. Automated pipelines should test safety, performance and cost impacts whenever an agent’s logic or model is updated, using multi-agent simulations to validate that new behaviours do not introduce regressions or unsafe interactions.
  • Hierarchical monitoring in multi-agent frameworks: Systems such as MegaAgent [128] decompose tasks into subtasks and dynamically form groups of agents. They employ hierarchical monitoring, where each agent tracks its actions with a checklist, group-level administrators oversee their agents, and system-level supervisors review outputs. This layered monitoring ensures that changes to one agent do not compromise the entire system.
  • Risk measurement and oversight policies: The NIST generative AI profile recommends documenting risks and updating measurement approaches regularly; it suggests using gradient-based attributions, occlusion and counterfactual prompts to improve explainability and transparency, and establishing policies for oversight across the lifecycle [126].
5.2.3 Application Security for LLM Agents
Security in AMAS systems must address vulnerabilities such as prompt injection, data exfiltration, and the misuse of external tools. According to OWASP [137, 138], prompt injection manipulates an LLM through malicious instructions, potentially causing the model to leak data, execute unauthorized commands or make harmful decisions. Multimodal models face cross-modal injection where malicious content in one modality influences the output of another. To mitigate these risks:
  • Prompt hygiene and hardening: Sanitize and filter user inputs, apply secure prefixes, and validate instructions before they reach an agent. Constraining the output format and employing guardrails can limit the model’s response surface [38].
  • Defense-in-depth design patterns: Existing techniques such as adversarial training and user confirmation provide partial protection, but research shows they do not guarantee safety. Design patterns [133] like Action-Selector and Plan-Then-Execute restrict an agent’s capabilities after processing untrusted input; the agent selects from a predefined list of allowed actions or produces a plan that cannot be altered by tool outputs, thereby isolating untrusted data from execution paths [58].
  • Least-privilege access and authentication: Enforce strong authentication for both human users and agents and apply role-based access controls so agents can only invoke the minimum set of tools necessary to perform their task. Continuous monitoring should detect anomalous requests; if an agent deviates from normal patterns, automated systems should flag or suspend it [152].
  • Post-processing and anomaly detection: Implement filters to detect and remove sensitive or inaccurate content from outputs and regularly retrain models to reduce hallucinations. Maintain robust logging and retention policies so that investigators can audit interactions without exposing private data.
@@ 5.2.3 Application Security for LLM Agents
Recent frameworks also emphasis hierarchical monitoring and cross-checking among agents to detect inconsistencies [128]. LangChain [153], AutoGen [98], and CrewAI [87] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise. Training agents with adversarial examples [134] and red-teaming [154] scenarios further strengthens robustness against malicious inputs.
5.2.4 Model Privacy and Data Protection
AMAS also poses unique privacy challenges because agents often share information via memory stores and message passing. Without strict controls, an agent may inadvertently leak sensitive data [40]. A layered privacy strategy should include:
  • Differential Privacy (DP): Adding calibrated noise to training data prevents individual data points from being re-identified and reduces the risk that models memorize or regurgitate personal information [139]. DP should be assessed on a case-by-case basis, since heavy noise can harm accuracy.
  • Data anonymization and minimization: Apply robust anonymization and pseudonymization techniques and regularly test them for effectiveness [140]. Limit data collection to what is strictly necessary for the agent’s task and enforce strict retention policies.
  • Secure Multi-Party Computation (MPC): MPC allows parties to jointly compute a function without revealing their private inputs. In privacy-preserving LLM services, secure MPC can protect user prompts from the model provider; for example, the PUMA framework uses MPC to ensure that prompts remain hidden during inference, while maintaining reasonable performance (it can evaluate a 7-billion-parameter model within minutes) [141].
  • Homomorphic Encryption (HE): HE enables computation on encrypted data and offers cryptographically provable privacy. Recent work demonstrates an encryption-friendly LLM architecture that uses LoRA fine-tuning and Gaussian kernels to speed up fully homomorphic encryption (FHE) operations, allowing encrypted inference with performance comparable to plaintext models [142].
  • Trusted Execution Environments (TEEs): TEEs provide hardware-enforced isolation that protects code and data [143]. They partition processors into secure and normal worlds; implementations include Intel SGX, AMD SEV and Arm TrustZone. In secure LLM inference, TEEs ensure that models run only inside authenticated enclaves, with attestation verifying model integrity.
  • Continuous monitoring and access controls: Encrypt data at rest and in transit, restrict access to authorized personnel, and use monitoring to detect data poisoning or unusual behaviour.
These techniques, combined with careful consent management and compliance with local regulations, help maintain confidentiality when agents share or process sensitive information.
Figure 6: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST AI RMF, ISO/IEC 42001, 42005, and OECD Principles. Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art. 13–14) [125] and the NIST AI RMF trustworthiness characteristic “Explainability & Interpretability” [3]; also supported by NISTIR 8312 (Four Principles of Explainable AI) [155]. Promotes well-justified (rationale) and human-interpretable outputs with traceable logs [119]. ModelOps: Anchored in the NIST AI RMF “Govern/Manage” functions [3] and ISO/IEC 42001:2023 (AI management system) [156]; secure SDLC controls via NIST SP 800-218/218A [^5] NIST, 2024. Emphasizes lifecycle gates, model versioning, rollback policies, and CI/CD with safety gates. Risk Management & Impact Assessment: EU AI Act (Art. 9, 72–73) [125], ISO/IEC 23894 (AI risk management) [^6] ISO/IEC 23894, and ISO/IEC 42005:2025 (AI impact assessment) [129]. Iterative hazard analysis, impact registers, incident response, and disclosure playbooks. Data Governance & Quality: EU AI Act (Art. 10) [125] mapped to the ISO/IEC 5259 series (data-quality model, measures, processes, governance) [^7] ISO/IEC 5259 series. Data minimization, lineage, coverage/bias checks, and sampling documentation. Robustness & Safety: EU AI Act (Art. 15) [125] and ISO/IEC 24029-1 (robustness assessment) [32]. Adversarial/perturbation testing, fault tolerance, fallback/kill-switch triggers, and rollback criteria. Application Security: Guided by OWASP LLM Top-10 [137] and OECD AI Principles on robustness [127]. Access controls, prompt/input mediation, sandboxing, egress controls, and auditable trails. Model Privacy: Aligned with GDPR (Art. 25) [34], HIPAA §164 [35], and OECD AI Principles [127]. Encryption in transit/at rest, DP budgeting, data minimization, and access-audit pass thresholds. Documentation & Traceability: EU AI Act (Arts. 11–12; Annex IV) [125]. Complete technical documentation, event logs, model/data cards, decision records, and regulator-ready evidence packages. Post-Market Monitoring: EU AI Act (Arts. 72–73) [125]. Telemetry for drift/misuse, continuous evaluations, feedback loops, controlled recalls, and stakeholder notifications. Governance: Grounded in GDPR, CCPA [144], and OECD AI Principles [127]. Institutional oversight, policy compliance, role-based accountability, regulatory traceability, and human-in-the-loop for high-risk/multi-agent deployments. Fairness & Human Rights: OECD AI Principles [127] and ISO/IEC TR 24027 (bias) [^8] ISO/IEC TR 24027. Lifecycle bias controls, affected-party impact reviews, and mitigation sign-offs.
5.2.5 Governance
Governance integrates ethical, legal and societal considerations into the design and operation of AI systems. Regulations such as the EU AI Act [125], NIST AI RMF [3] and OECD AI [127] provide guidance that aligns with TRiSM pillars of Explainability, ModelOps, Security and Privacy.
  • EU AI Act - Human Oversight (Article 14): High-risk AI systems must be designed so that natural persons can effectively oversee them and prevent or minimize risks to health, safety or fundamental rights [1]. The AI system should enable overseers to understand capabilities and limitations, detect anomalies, interpret outputs, decide not to use the system, or stop its operation through a ‘stop’ button. For certain identification tasks, decisions must be independently verified by at least two competent individuals.
  • OECD AI Principles: These intergovernmental principles, updated in 2024, promote human-centric, trustworthy AI [127]. They emphasise transparency and explainability: AI actors should provide meaningful information about data sources, factors and logic to allow affected individuals to understand and challenge outputs. Robustness, security and safety require AI systems to function appropriately under normal and adverse conditions and include mechanisms to override or decommission systems to prevent harm. Accountability means AI actors are responsible for the proper functioning of AI systems; they must ensure traceability of datasets and decisions and apply systematic risk management throughout the lifecycle.
  • NIST AI Risk Management Framework (AI RMF): The AI RMF and its Generative AI profile outline four core functions: govern, map, measure and manage [3]. For generative AI, NIST recommends documenting risks and potential impacts, applying interpretability techniques (gradient-based attributions, occlusion, counterfactual prompts) to improve transparency, and establishing policies for oversight across the AI lifecycle.
  • Auditability and role-based access: Agentic AI systems should maintain auditable logs of agent actions, decisions and data flows and assign clear roles (e.g., reasoning, verification, summarisation) so that responsibilities are transparent [157]. Role-based access controls help enforce accountability and prevent unauthorized agents from accessing sensitive functions.
  • Global harmonisation: Many jurisdictions (e.g., GDPR, CCPA) require data subject rights, consent and privacy by design. Frameworks like the OECD Principles and NIST AI RMF promote international cooperation and interoperability.
These governance frameworks (as illustrated in Figure 6) underscore the need for transparent accountability structures, human oversight, robust privacy controls, and documented processes in AMAS. TRiSM provides a practical scaffolding for operationalizing these requirements, aligning lifecycle management, security, privacy and governance to build trustworthy Agentic AI.
Next, we discuss the evaluations in AMAS.